Circular No. 8600

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2005ea IN MCG +10-16-61 D. Lane and P. Gray, Stillwater Lake, NS, Canada, report their dis-

covery of an apparent supernova (mag ∼ 17.0) on an unfiltered CCD image (cf. IAUC 8462) taken on Sept. 6.016 UT, with a confirming image showing the new object at mag ∼ 16.3 on Sept. 13.015. SN 2005ea is located at α = 11^{h}06^{m}47^{s}.89, δ = +57^{o}41′07′′.5 (equinox 2000.0), which is ∼ 14′′ west of the nucleus of the galaxy MCG +10-16-61. Nothing is visible at this location on their images taken on Apr. 22 (limiting mag ∼ 17.7), June 6 (limiting mag ∼ 18.5), and Aug. 18 (limiting mag ∼ 17.8), and nothing is present on Digitized Sky Survey images from 1994 (limiting red mag ∼ 21.0) and 1988 (limiting blue mag ∼ 20.5). A. Sehgal, Osoyoos, BC, Canada, reports that 2005ea appeared at mag 17.2 and at position end figures 47^{s}.89, 07′′.4 on an unfiltered CCD image taken with a 0.50-m reflector on Sept. 13.12.

V1647 ORIONIS AND ASSOCIATED NEBULA C. Aspin, Gemini Observatory; and B. Reipurth, University of Hawaii,

report on observations made at the Gemini 8-m telescope on Mauna Kea. GMOS optical images obtained on Aug. 30 UT show that V1647 Ori (cf. IAUC 8354), upon its re-emergence in the morning twilight, has faded by ≈ 1 magnitude in the r′ passband since Jan. 8. McNeil’s nebula itself (cf. IAUC 8284) has also faded considerably. The duration of the outburst so far exceeds 21 months. Optical spectra still show strong Hα emission with blue-shifted absorption. Both the depth and extent of the strong absorption component are less than previously observed, suggesting that the powerful wind is weakening.

COMET 169P/NEAT M. Jäger, Vienna, reports that his co-added CCD images of this comet

(cf. IAUC 8578, 8591) taken on Sept. 6.11 and 7.11 UT with a 0.20-m Schmidt telescope, with the comet low in the morning sky, show a coma of diameter 2′-2′.5 and (on Sept. 7.11) a faint tail ≈ 10′ long. A. Hale, Cloudcroft, NM, reports that his visual observations on Sept. 12.48 (0.2-m reflector) and 13.47 (0.41-m reflector) show the diffuse comet at total mag 11.5 with a coma diameter of 1′.5 and some slight condensation. Hale had previously observed the comet in early August, when it appeared as an essentially stellar object of mag ≈ 14.5–15.0.

2005 September 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green